OpenStudy (anonymous):

How do you calculate the density of a gas, based on its temperature and pressure? I've done it before, but I need a good way of remembering the process.

OpenStudy (anonymous):

Its just ideal gas equation If PV=nRT Now n=m/M (Or mass by molecular mass) substitue to get PM=m/V RT Now m/V is nothing but density if the gas PM= dRT

OpenStudy (anonymous):

u can use the ideal gas law V.P=n.R.T Since: n=mass of gas/molecular mass (M.M) Therefore V.P=( m/MM).R.T Therefore: P=(m/M.M.V).R.T since: d=m/v so P=(d.R.T)/M.M so d=(P. M.M)/RT To find the density of a gas, take the P in atm times the molar mass, then divide by R (0.08206 L atm / K mol) and T (in kelvins)

OpenStudy (anonymous):

Felix: Did you get it?